OpenStudy (anonymous):

if (a/x)-(b/y)=c and xy=1/c, then bx= (a)1-ay (b)ay (c)ay+1 (d)ay-1 (e)2ay

OpenStudy (anonymous):

\[\frac{ a }{ x }-\frac{ b }{ y }=c,xy=\frac{ 1 }{ c }\] \[\frac{ ay-bx}{ xy }=c\] \[ay-bx=cxy=c \frac{ 1 }{ c }=1\] bx=ay-1